The revenue from selling x units of a product is given by
y = -0.0002x^2 + 20x. How many units must be sold in
order to have the greatest revenue? (Find the x-coordinate
of the vertex of the parabola.)

Answer by stanbon(75887)

The revenue from selling x units of a product is given by
y = -0.0002x^2 + 20x. How many units must be sold in
order to have the greatest revenue? (Find the x-coordinate
of the vertex of the parabola.)
-----
maximum revenue occurs when x = -b/(2a) = -20/(2*-0.0002) = 50,000
-------------------------
Cheers,
Stan H.